two steamers leave a port on a river at the same time, traveling in opposite direction. each is traveling at 22 mph.how long will them to be 110 mi apart?
Answer by mananth(16946)

speed of ships = 22 mph
opposite direction away from each other
combined speed = 22+22=44mph
distance = 110 miles
t=d/r
t=110/44
t= 2.5 hours